A Square Plate of side l has mass per unit area mew find its moment of inertia about an Axis passing through the centre and perpendicular to its plane

Grade:11

1 Answers

Arun
6 years ago
Dear Gaurav
 
Moment of Inertia of a square plate of side ‘l’ and mass ‘m’ about an axis passing through its centre and perpendicular to its plane is, I = (1/6)ml²
 
Length of the diagonal is  l\sqrt2.
So the distance of the corner from the centre is
= l \sqrt2/2
So the moment of inertia of plate about an axis passing through centre is
I = (1/6)ML² + M(L\sqrt2/2)²
I = (1/6)ML² + (1/2) ML² = (2/3) ML²
 
Regards
Arun (askIITians forum expert)